OpenStudy (anonymous):

Write the equation of a line in slope-intercept form going through the points (3,-4) and (1,2)

OpenStudy (anonymous):

Slope is -3 you get this from m = (2-(-4))/(1-3) (y-y1)=m(x-x1) (y-2)=(-3)(x-1) y=-3(x-1)+2

OpenStudy (anonymous):

y=-3(x-1)+2 distribute the -3 y=-3x+3+2 y=-3x+5 is the equation

OpenStudy (anonymous):

Thanks for the help!